# PicoAudio: Enabling Precise Timestamp and Frequency Controllability of Audio Events in Text-to-audio Generation

### v1

@@ Abstract
Recently, audio generation tasks have attracted considerable research interests. Precise temporal controllability is essential to integrate audio generation with real applications. In this work, we propose a temporal controlled audio generation framework, PicoAudio. PicoAudio integrates temporal information to guide audio generation through tailored model design. It leverages data crawling, segmentation, filtering, and simulation of fine-grained temporally-aligned audio-text data. Both subjective and objective evaluations demonstrate that PicoAudio dramantically surpasses current state-of-the-art generation models in terms of timestamp and occurrence frequency controllability. The generated samples are available on the demo website https://PicoAudio.github.io.
@@ 1 Introduction
In this work, we propose PicoAudio which enables Precise tImestamp and frequency COntrollability of audio events, by leveraging data simulation [^1] Simulated datasets for training and evaluation are available at https://github.com/PicoAudio/PicoAudio, tailored model designs, and preprocessing with large language model. We focus on timestamp and frequency control, while other temporal conditions (e.g., ordering and interval) can be converted into timestamps through textual reasoning, akin to transforming frequency into timestamps in our experiment. PicoAudio proposes a pipeline to simulate data with temporally-aligned annotations. The pipeline entails crawling data from the Internet, segmenting and filtering audio clips to gather high-quality audio segments, as well as simulating to synthesize realistic audio. PicoAudio introduces tailored modules for temporal control. (a) Timestamp control is accomplished by incorporating customized input, namely timestamp caption. With the assistance of large language model (LLM) [15], (b) frequency control, (c) ordering via multi-event timestamp control and (d) multi-event frequency control can be implemented, as shown in Figure 1. Beyond (a)-(d), PicoAudio can achieve arbitrary precise temporal control as long as the LLM is capable of converting the requirement into timestamp captions, which is straightforward for LLM when prompted with simulated data. Our contributions encompass the following:

### v2

@@ Abstract
Recently, audio generation tasks have attracted considerable research interests. Precise temporal controllability is essential to integrate audio generation with real applications. In this work, we propose a temporal controlled audio generation framework, PicoAudio. PicoAudio integrates temporal information to guide audio generation through tailored model design. It leverages data crawling, segmentation, filtering, and simulation of fine-grained temporally-aligned audio-text data. Both subjective and objective evaluations demonstrate that PicoAudio dramantically surpasses current state-of-the-art generation models in terms of timestamp and occurrence frequency controllability. The generated samples are available on the demo website https://zeyuxie29.github.io/PicoAudio.github.io.
@@ 1 Introduction
In this work, we propose PicoAudio which enables Precise tImestamp and frequency COntrollability of audio events, by leveraging data simulation [^1] Simulated datasets for training and evaluation are available at https://github.com/zeyuxie29/PicoAudio, tailored model designs, and preprocessing with large language model. We focus on timestamp and frequency control, while other temporal conditions (e.g., ordering and interval) can be converted into timestamps through textual reasoning, akin to transforming frequency into timestamps in our experiment. PicoAudio proposes a pipeline to simulate data with temporally-aligned annotations. The pipeline entails crawling data from the Internet, segmenting and filtering audio clips to gather high-quality audio segments, as well as simulating to synthesize realistic audio. PicoAudio introduces tailored modules for temporal control. (a) Timestamp control is accomplished by incorporating customized input, namely timestamp caption. With the assistance of large language model (LLM) [15], (b) frequency control, (c) ordering via multi-event timestamp control and (d) multi-event frequency control can be implemented, as shown in Figure 1. Beyond (a)-(d), PicoAudio can achieve arbitrary precise temporal control as long as the LLM is capable of converting the requirement into timestamp captions, which is straightforward for LLM when prompted with simulated data. Our contributions encompass the following: